# Mario Shannon

### v1

@@ 1.2. Graph of Anosov flows
Let $A\in\mathrm{SL}(2,\mathbb{Z})$ be a matrix with $\textsc{tr}(A)\geq 3$ satisfying that $A$ is conjugated to $A^{-1}$ in $\mathrm{GL}(2,\mathbb{Z})$. Then, given any periodic orbit $\gamma$ of the suspension flow $(\phi_{t}^{A},M_{A})$ there exists $m_{0}=m_{0}(\gamma)>0$ such that, for every $m\geq m_{0}$ there is a set $\mathcal{Q}(\gamma,m)$ containing infinitely many $\phi_{t}^{A}$-periodic orbits, satisfying that
$(\phi_{t}^{A},M_{A})\xrightarrow[]{(\gamma,m),(\alpha,-m)}(\phi_{t}^{A},M_{A}),\ \forall\ \alpha\in\mathcal{Q}(m,\gamma).$
The image of the periodic orbit $\alpha$ under this surgery equals the periodic orbit $\gamma$, up to a self-orbital equivalence of $(\phi_{t}^{A},M_{A})$.
In other words, for $A\simeq A^{-1}$ we find infinitely many different loops of length=2 in $\mathbf{G}_{\circ}$ based at $(\phi_{t}^{A},M_{A})$. Each of these loops consists in making one surgery of parameter $m$ in one periodic orbit $\gamma$, together with one surgery of parameter $-m$ in a different periodic orbit $\alpha$, where $m\in\mathbb{Z}$ may be taken arbitrarily big. It is important to remark that for every fixed $m\in\mathbb{Z}$, this set of loops is small, in the following sense:
@@
Let $\mathbb{P}$ be the set of periodic orbits of $(\phi_{t}^{A},M_{A})$, and for some fixed $m>0$ define $\mathcal{Q}^{2}(m)=\{(\gamma_{1},\gamma_{2})\in\mathbb{P}^{2}:\gamma_{2}\in\mathcal{Q}(\gamma_{1},m)\}.$ Denote by $\mathbb{P}^{2}_{t}$ and $\mathcal{Q}^{2}_{t}(m)$ the respective subsets formed by pairs $(\gamma_{1},\gamma_{2})$ with $\textsl{per}(\gamma_{i})\leq t$. Then, it is verified that
$\lim_{t\to+\infty}\frac{|\mathcal{Q}^{2}_{t}(m)|}{|\mathbb{P}^{2}_{t}|}=0.$
Observe that the loops described in the theorem above are not, in principle, all the closed loops of length $=2$ based at $(\phi_{t}^{A},M_{A})$.
Question 7.
Is it possible to estimate the growth rate as a function of the period, of the total number of loops in $\mathbf{G}_{\circ}$ of length $=2$ (or any other fixed length) based at $(\phi_{t}^{A},M_{A})$
@@
As a final remark, when the matrix $A$ is not conjugated to $A^{-1}$ in $\mathrm{GL}(2,\mathbb{Z})$, the construction of Theorem A provides paths of length=2 connecting two suspensions
$(\phi_{t}^{A},M_{A})\xrightarrow[]{(\gamma,m),(\alpha,-m)}(\phi_{t}^{B},M_{B}),\ \forall\ \alpha\in\mathcal{Q}(m,\gamma),$
where $B$ is conjugated to either $A$ or $A^{-1}$. We do not know, however, how to determine exactly the conjugacy class of $B$ in this case.
@@ 2. $\infty$-many closed paths
The construction of the sets $\mathcal{Q}(\gamma,m)$ in Theorem A is an elementary application of three results due to Plante, Thurston and Barthelme-Fenley. We will summarize these results first and explain how to construct the paths stated in Theorem A at the end.
@@ 2.1. Anosov flows on sol-manifolds
A suspension Anosov flow is one constructed by choosing a hyperbolic matrix $A\in\mathrm{GL}(2,\mathbb{Z})$ and making the quotient of the space $\mathbb{T}^{2}\times\mathbb{R}$ by the action $(x,t)\mapsto(A(x),t-1)$, where $\mathbb{T}^{2}=\mathbb{R}^{2}/\mathbb{Z}^{2}$. This gives a closed $3$-manifold $M_{A}$ equipped with an Anosov flow $\phi_{t}^{A}:M_{A}\to M_{A}$ induced from the action $\tilde{\phi}_{t}(x,s)=(x,s+t)$ on $\mathbb{T}^{2}\times\mathbb{R}$. The stable and unstable invariant bundles of the flow are induced from the contracting and expanding eigenspaces of the matrix $A$ acting on $\mathbb{T}^{2}$.
@@
Let $\phi_{t}:M\to M$ be an Anosov flow on a closed 3-manifold with solvable fundamental group. Then, $(\phi_{t},M)$ is orbitally equivalent to a suspension Anosov flow $(\phi_{t}^{A},M_{A})$, where $A\in\mathrm{GL}(2,\mathbb{Z})$ is a hyperbolic matrix.
On the 3-manifold $M_{A}$ there is another Anosov flow $\bar{\phi}_{t}^{A}:M_{A}\to M_{A}$ and another torus bundle $\pi_{-}:M_{A}\to\mathbb{R}/\mathbb{Z}$ with monodromy $A^{-1}$, such that its fibers $\pi_{-}^{-1}(\theta)$ are global transverse sections for $\bar{\phi}_{t}^{A}$ with first return conjugated to the monodromy. These are obtained by reversing the time direction of $\phi_{t}^{A}$ and reversing the orientation of the base in the bundle projection $\pi_{+}$. Both flows $\phi_{t}^{A}$ and $\bar{\phi}_{t}^{A}$ define the same foliation by orbits of $M_{A}$, up to changing the orientation of the orbits, and both maps $\pi_{+}$ and $\pi_{-}$ define the same fibration by tori of $M_{A}$, up to changing the coorientation of the fibers. There is a rigidity property of these foliations that allows to sharpen the statement in theorem 9 above:
@@
This folkloric statement follows by direct application of Thurston’s fibered faces theory, see e.g. [18]. Since $H^{1}(M_{A},\mathbb{Z})$ has rank $=1$ then there is a unique fibration of $M_{A}$ by closed surfaces, up to isotopy. As a consequence one has that:
@@
Let $A\in\mathrm{GL}(2,\mathbb{Z})$ be a hyperbolic matrix and let $M_{A}$ be the 3-manifold obtained by suspension of $A$. Then every Anosov flow $\psi_{t}:M_{A}\to M_{A}$ is orbitally equivalent to either $\phi_{t}^{A}$ or $\bar{\phi}_{t}^{A}$, under an orbital equivalence isotopic to the identity on $M_{A}$.
@@
If $\psi_{t}:M_{A}\to M_{A}$ is any Anosov flow, since $\pi_{1}(M_{A},x_{0})$ is solvable then by theorem 9 it is equivalent to a suspension flow $(\phi_{t}^{B},M_{B})$ with $B\in\mathrm{GL}(2,\mathbb{Z})$ hyperbolic. Hence, there is a fibration $\pi:M_{A}\to\mathbb{R}/\mathbb{Z}$ by tori $\pi^{-1}(\theta)$, $\theta\in\mathbb{R}/\mathbb{Z}$ each of them being a global transverse section for $\psi_{t}$ with monodromy conjugated to $B$. By theorem 10 there is a homeomorphism $H:M_{A}\to M_{A}$ isotopic to the identity such that $\pi=\pi_{\varepsilon}\circ H$, where $\varepsilon=+,-$. Hence, the monodromy $B$ of the bundle $\pi$ is conjugated either $A$ of $A^{-1}$, and the homeomorphism $H$ realizes an orbital equivalence between $(\psi_{t},M_{A})$ and either $(\phi_{t}^{A},M_{A})$ or $(\bar{\phi}_{t}^{A},M_{A})$. ∎
Finally, we will use a property of the foliation by orbits of $(\phi_{t}^{A},M_{A})$, stated in proposition 12 below. Let’s introduce some terminology.
@@
Let $(\phi_{t}^{A},M_{A})$ be the suspension Anosov flow generated by a hyperbolic matrix $A\in\mathrm{GL}(2,\mathbb{Z})$, and let $\gamma$ and $\delta$ be two periodic orbits. If there exists a homeomorphism $M_{A}\setminus\gamma\to M_{A}\setminus\delta$ preserving the orientations at infinity of $\gamma$ and $\delta$, then there exists a self-orbital equivalence $\bar{H}:(\phi_{t}^{A},M_{A})\to(\phi_{t}^{A},M_{A})$ such that $\bar{H}(\gamma)=\delta$.
@@
To show the proposition it is enough to show the existence of a homeomorphism $H:M_{A}\to M_{A}$ such that $H(\gamma)=\delta$, preserving the orientations of these curves. To see this, let $\psi_{t}$ be the flow on $M_{A}$ obtained by pushing forward the action of $\phi_{t}^{A}$ with $H$. Then, by proposition 11 there is an orbital equivalence  [^1] Note that $\psi_{t}$ is technically a topological Anosov flow. However, proposition 11 is still valid due to the results in [16]. See also [9]. $H^{*}:(\psi_{t},M_{A})\to(\phi_{t}^{A},M_{A})$ isotopic to the identity. Since $H^{*}(\delta)$ is isotopic to $\delta$ then $H^{*}(\delta)=\delta$, because for every suspension Anosov flow there exist at most one periodic orbit per (free) isotopy class (see e.g. [2]). Hence, the homeomorphism $\bar{H}=H^{*}\circ H$ is a self-orbital equivalence of the suspension flow taking $\bar{H}:\gamma\mapsto\delta$.
@@
Since $[\delta]\neq 0$ for every periodic orbit $\delta$ of the suspension flow, we conclude that $m=0$. Therefore, using the claim above we can modify the homeomorphism $h$ inside the tubular neighborhoods $W_{\gamma}$ and $W_{\delta}$, and obtain a globally defined $H:M_{A}\to M_{A}$ as required. ∎
2.2. Hyperbolic Dehn surgery.
Let $M$ be a closed 3-manifold, $\gamma$ a simple closed curve, and assume that $M\setminus\gamma$ admits a complete hyperbolic metric. A classical theorem of Thurston (see [19]) states that making a Dehn surgery on $M$ along $\gamma$ will produce a hyperbolic closed 3-manifold, for all except a finite number of surgery slopes. This method for constructing hyperbolic 3-manifolds is known as hyperbolic Dehn surgery.
Hyperbolic Dehn surgery can be implemented in the context of Anosov flows and Fried surgeries, to produce Anosov flows on hyperbolic 3-manifolds.  [^2] This method goes back to Goodman in [10], where another version of the surgery is used. See [15] for an explanation of how the Goodman and Fried surgeries are related. For instance, in the case of an Anosov flow equivalent to the suspension of a hyperbolic matrix with positive eigenvalues, it can be shown that the complement of every periodic orbit admits a complete hyperbolic metric, by showing that it is atoroidal. This can be accomplished by showing (as in the case of proposition 11) that the manifold obtained by removing the periodic orbit has only one class of incompressible embedded surfaces, namely, the torus fibers of the suspension but punctured along the removed orbit.
We get the following statement:
@@
Consider the suspension flow $(\phi_{t}^{A},M_{A})$ generated by $A\in\mathrm{SL}(2,\mathbb{Z})$ with $\textsc{tr}(A)\geq 3$. Then, for every $\phi_{t}^{A}$-periodic orbit $\gamma$ the Fried surgery along $\gamma$ with slope $m$ will produce an Anosov flow on a hyperbolic 3-manifold, for all except a finite number of $m\in\mathbb{Z}$.
2.3. $\mathbb{R}$-covered Anosov flows on hyperbolic 3-manifolds.
Let $\phi_{t}:M\to M$ be an Anosov flow on a closed 3-manifold and denote by $\mathcal{O}$ the corresponding foliation of $M$ by $\phi_{t}$-orbits. Let $\tilde{M}$ be the universal cover of $M$ and let $\tilde{\mathcal{O}}$ be the lift of the foliation $\mathcal{O}$. The following is a fundamental result due to Barbot and Fenley:

### v2

@@ 1.3. Infinitely many loops of length 2 based at a suspension
Let $A\in\mathrm{SL}(2,\mathbb{Z})$ be a matrix with $\operatorname{\mathrm{tr}}(A)\geq 3$ satisfying that $A$ is conjugated to $A^{-1}$ in $\mathrm{GL}(2,\mathbb{Z})$. Then, given any periodic orbit $\gamma$ of the suspension flow $(\phi^{A},M_{A})$ there exists $m_{0}=m_{0}(\gamma)>0$ such that, for every integer $m$ with $|m|\geq m_{0}$, there is a set $\mathcal{Q}(\gamma,m)$ containing infinitely many $\phi^{A}$-periodic orbits, satisfying that
$(\phi^{A},M_{A})]\xrightarrow[]{(\gamma,m),(\alpha,-m)}[(\phi^{A},M_{A})],\ \text{for every}\ \alpha\in\mathcal{Q}(m,\gamma).$
Moreover, let $\Pi:\textrm{Per}(\phi^{A},M_{A})\to\textrm{Per}(\phi^{A},M_{A})$ be the correspondence that identifies periodic orbits before and after this surgery operation. Then, the image of $\alpha$ under the correspondence $\Pi$ equals the periodic orbit $\gamma$, up to a self-orbital equivalence of $(\phi^{A},M_{A})$.
In other words, for hyperbolic matrices $A$ conjugated to its inverse, we find infinitely many different loops of length $2$ in $\mathbf{G}_{\circ}$ based at the class $(\phi^{A},M_{A})]$. Each of these loops consists in making one surgery of slope $m$ along one periodic orbit $\gamma$, together with one surgery of slope m-m along a different periodic orbit $\alpha$ chosen from the set $\mathcal{Q}(m,\gamma)$, and $m\in\mathbb{Z}$ may be taken arbitrarily big. Since $\mathcal{Q}(m,\gamma)$ is infinite, it contains orbits $\alpha$ which cannot be sent into $\gamma$ using a self-orbit equivalence of $(\phi^{A},M_{A})$. The fact that $\Pi(\alpha)=\gamma$ up to self-orbital equivalence means that the way in which the periodic orbits of $(\phi^{A},M_{A})$ identify with themselves under this double-surgery operation is not trivial, in the sense that $\Pi$ is not induced from a self-orbital equivalence of $(\phi^{A},M_{A})$. This means that the surgery encoded on this loop is not trivial either.
It is important to remark that, for every fixed $m$ in $\mathbb{Z}$ and $\gamma$ in $\textrm{Per}(\phi^{A},M_{A})$, the set of loops described in Theorem A is small, in the following sense:
@@
Let $\gamma$ be a periodic orbit of $(\phi^{A},M_{A})$, let $m$ be an integer satisfying that $|m|\geq m_{0}(\gamma)$, and let $\mathcal{Q}(\gamma,m)$ be the set of periodic orbits described in Theorem A. Then:
$\lim_{t\to\infty}\frac{\left|\left\{\alpha\in\mathcal{Q}(\gamma,m):\operatorname{\mathrm{per}}(\alpha)<t\right\}\right|}{\left|\left\{\alpha\in\textrm{Per}(\phi^{A},M_{A})):\operatorname{\mathrm{per}}(\alpha)<t\right\}\right|}=0.$
Observe that the previous asymptotic is given in terms of a fixed $m$ and $\gamma$. It would be interesting to find an asymptotic on the total number of loops
$(\phi^{A},M_{A})]\xrightarrow[]{(\gamma,m),(\alpha,-m)}[(\phi^{A},M_{A})],\ \alpha\in\mathcal{Q}(\gamma,m)$
with $m$ and $\gamma$ varying among all possible choices.
Question 9.
Is it possible to estimate the growth rate, as a function of the period, of the total number of loops in $\mathbf{G}_{\circ}$ of length 2 (or any other fixed length) based at $(\phi^{A},M_{A})]$ ?
@@
As a final remark, when the matrix $A$ is not conjugated to $A^{-1}$ in $\mathrm{GL}(2,\mathbb{Z})$, the construction of Theorem A provides paths of length $2$ connecting two suspensions
$(\phi^{A},M_{A})]\xrightarrow[]{(\gamma,m),(\alpha,-m)}[(\phi^{B},M_{B})],\ \forall\ \alpha\in\mathcal{Q}(m,\gamma),$
where $B$ is conjugated to either $A$ or $A^{-1}$. However, determining the exact conjugacy class of $B$ in this case is more difficult.
@@ 2. $\infty$-many closed paths
The construction of the sets $\mathcal{Q}(\gamma,m)$ in Theorem A is an elementary application of three results due to Plante, Thurston and Barthelmé-Fenley. We will summarize these results first and explain how to construct the paths stated in Theorem A at the end.
@@ 2.1. Anosov flows on sol-manifolds
A suspension Anosov flow is one constructed by choosing a hyperbolic matrix $A\in\mathrm{GL}(2,\mathbb{Z})$ and making the quotient of the space $\mathbb{T}^{2}\times\mathbb{R}$ by the action $(x,t)\mapsto(A(x),t-1)$, where $\mathbb{T}^{2}=\mathbb{R}^{2}/\mathbb{Z}^{2}$. This gives a closed $3$-manifold $M_{A}$ equipped with an Anosov flow $\{\phi_{t}^{A}:M_{A}\to M_{A}\}_{t\in\mathbb{R}}$ induced from the action $\tilde{\phi}_{t}(x,s)=(x,s+t)$ on $\mathbb{T}^{2}\times\mathbb{R}$. The stable and unstable invariant bundles of the flow are induced from the contracting and expanding eigenspaces of the matrix $A$ acting on $\mathbb{T}^{2}$.
@@
Let $\{\phi_{t}:M\to M\}_{t\in\mathbb{R}}$ be an Anosov flow on a closed 3-manifold with solvable fundamental group. Then, $(\phi,M)$ is orbitally equivalent to a suspension Anosov flow $(\phi^{A},M_{A})$, where $A\in\mathrm{GL}(2,\mathbb{Z})$ is a hyperbolic matrix.
On the 3-manifold $M_{A}$ there is another Anosov flow $\{\bar{\phi}_{t}^{A}:M_{A}\to M_{A}\}_{t\in\mathbb{R}}$ and another torus bundle $\pi_{-}:M_{A}\to\mathbb{R}/\mathbb{Z}$ with monodromy $A^{-1}$, such that its fibers $\pi_{-}^{-1}(\theta)$ are global transverse sections for $\bar{\phi}^{A}$ with first return conjugated to the monodromy. These are obtained by reversing the time direction of $\phi^{A}$ and reversing the orientation of the base in the bundle projection $\pi_{+}$. Both flows $\phi^{A}$ and $\bar{\phi}^{A}$ define the same foliation by orbits of $M_{A}$, up to changing the orientation of the orbits, and both maps $\pi_{+}$ and $\pi_{-}$ define the same fibration by tori of $M_{A}$, up to changing the coorientation of the fibers. There is a rigidity property of these foliations that allows to sharpen the statement in Theorem 11 above:
@@
This folkloric statement follows by direct application of Thurston’s fibered faces theory, see e.g. [26]. Since $H^{1}(M_{A},\mathbb{Z})$ has rank equal to $1$ then there is a unique fibration of $M_{A}$ by closed surfaces, up to isotopy. As a consequence one has that:
@@
Let $A\in\mathrm{GL}(2,\mathbb{Z})$ be a hyperbolic matrix and let $M_{A}$ be the 3-manifold obtained by suspension of $A$. Then every Anosov flow $\{\psi_{t}:M_{A}\to M_{A}\}_{t\in\mathbb{R}}$ is orbitally equivalent to either $\phi^{A}$ or $\bar{\phi}^{A}$, under an orbital equivalence isotopic to the identity on $M_{A}$.
@@
If $\{\psi_{t}:M_{A}\to M_{A}\}_{t\in\mathbb{R}}$ is any Anosov flow, since $\pi_{1}(M_{A},x_{0})$ is solvable then by Theorem 11 it is equivalent to a suspension flow $(\phi^{B},M_{B})$ with $B\in\mathrm{GL}(2,\mathbb{Z})$ hyperbolic. Hence, there is a fibration $\pi:M_{A}\to\mathbb{R}/\mathbb{Z}$ by tori $\pi^{-1}(\theta)$, $\theta\in\mathbb{R}/\mathbb{Z}$ each of them being a global transverse section for $\psi$ with monodromy conjugated to $B$. By Theorem 12 there is a homeomorphism $H:M_{A}\to M_{A}$ isotopic to the identity such that $\pi=\pi_{\varepsilon}\circ H$, where $\varepsilon=+,$-. Hence, the monodromy $B$ of the bundle $\pi$ is conjugated either $A$ of $A^{-1}$, and the homeomorphism $H$ realizes an orbital equivalence between $(\psi,M_{A})$ and either $(\phi^{A},M_{A})$ or $(\bar{\phi}^{A},M_{A})$. ∎
Finally, we will use a property of the foliation by orbits of $(\phi^{A},M_{A})$, stated in Proposition 14 below. Let’s introduce some terminology.
@@
Let $(\phi^{A},M_{A})$ be the suspension Anosov flow generated by a hyperbolic matrix $A\in\mathrm{GL}(2,\mathbb{Z})$, and let $\gamma$ and $\delta$ be two periodic orbits. If there exists a homeomorphism $M_{A}\setminus\gamma\to M_{A}\setminus\delta$ preserving the orientations at infinity of $\gamma$ and $\delta$, then there exists a self-orbital equivalence $\bar{H}:(\phi^{A},M_{A})\to(\phi^{A},M_{A})$ such that $\bar{H}(\gamma)=\delta$.
@@
To show the proposition it is enough to show the existence of a homeomorphism $H:M_{A}\to M_{A}$ such that $H(\gamma)=\delta$, preserving the orientations of these curves. To see this, let $\psi$ be the flow on $M_{A}$ obtained by pushing forward the orbits of $\phi^{A}$ with $H$. Then, by Proposition 13 there is an orbital equivalence  [^2] Note that $\psi$ is technically a topological Anosov flow. However, Proposition 13 is still valid due to the results in [24]. See also [9]. $H^{*}:(\psi,M_{A})\to(\phi^{A},M_{A})$ isotopic to the identity. Since $H^{*}(\delta)$ is isotopic to $\delta$ then $H^{*}(\delta)=\delta$, because for every suspension Anosov flow there exist at most one periodic orbit per (free) isotopy class (see e.g. [2]). Hence, the homeomorphism $\bar{H}=H^{*}\circ H$ is a self-orbital equivalence of the suspension flow taking $\bar{H}:\gamma\mapsto\delta$.
@@
Since $\delta]\neq 0$ for every periodic orbit $\delta$ of the suspension flow, we conclude that $m=0$. Therefore, using the claim above we can modify the homeomorphism $h$ inside the tubular neighborhoods $W_{\gamma}$ and $W_{\delta}$, and obtain a globally defined $H:M_{A}\to M_{A}$ as required. ∎
2.2. Hyperbolic Dehn surgery
Let $M$ be a closed 3-manifold, $\gamma$ a simple closed curve, and assume that $M\setminus\gamma$ admits a complete hyperbolic metric. A classical theorem of Thurston (see [27], [22]) states that making a Dehn surgery on $M$ along $\gamma$ will produce a hyperbolic closed 3-manifold, for all except a finite number of surgery slopes. This method for constructing hyperbolic 3-manifolds is known as hyperbolic Dehn surgery.
Hyperbolic Dehn surgery can be implemented in the context of Anosov flows and Fried surgeries to produce Anosov flows on hyperbolic 3-manifolds. This method goes back to Goodman [^3] In Goodman’s work [10] other version of the Dehn surgery is used. See [23] for an explanation of how the Goodman and Fried surgeries are related. in [10], where the first examples of Anosov flows on hyperbolic 3-manifolds were obtained. These examples can be described as follows: take the suspension flow $(\phi^{A},M_{A})$ associated to the matrix $A=[2,1,1,1]$ and make a Fried surgery of slope $m\neq 0$ along the periodic orbit $\gamma_{0}$ associated to the (unique) fixed point of $A$. It is a well-known fact (cf. Chapter 8 of [13]) that the complement of $\gamma_{0}$ in $M_{A}$ is homeomorphic to the complement of the figure-eight knot in the sphere $\mathbb{S}^{3}$, which admits a complete hyperbolic metric (cf. [27]). Hence, when the slope $m$ of the surgery is chosen big enough, we obtain an Anosov flow on a hyperbolic 3-manifold. We refer to [29] for a treatment of the Anosov flows obtained in this particular example and its relation to the figure-eight knot complement.
The method described above is completely general and does not depend on the special matrix $A=[2,1,1,1]$. It can be shown that, for every $A\in\mathrm{SL}(2,\mathbb{Z})$ with $\operatorname{\mathrm{tr}}(A)\geq 3$, the complement of any periodic orbit $\gamma$ of the suspension Anosov flow $(\phi^{A},M_{A})$ admits a complete hyperbolic metric. This follows from Thurston’s hyperbolization theorem at [28]. Since $M_{A}\setminus\gamma$ fibers over the circle with fibers homeomorphic a punctured torus, and since the monodromy $A$ on the fibers is of pseudo-Anosov type, then case (iii) of Theorem 0.1 at [28] applies and $M_{A}\setminus\gamma$ is hyperbolic. We refer to [12] for a general treatment of the problem of whether the complement of a periodic orbit of an Anosov flow admits a complete hyperbolic metric. See also [16], [17].
Putting all together, we get the following statement:
@@ 2.2. Hyperbolic Dehn surgery
Consider the suspension flow $(\phi^{A},M_{A})$ generated by $A\in\mathrm{SL}(2,\mathbb{Z})$ with $\operatorname{\mathrm{tr}}(A)\geq 3$. Then, for every $\phi^{A}$-periodic orbit $\gamma$ the Fried surgery along $\gamma$ with slope $m$ will produce an Anosov flow on a hyperbolic 3-manifold, for all except a finite number of $m\in\mathbb{Z}$.
2.3. $\mathbb{R}$-covered Anosov flows on hyperbolic 3-manifolds
Let $\{\phi_{t}:M\to M\}_{t\in\mathbb{R}}$ be an Anosov flow on a closed 3-manifold and denote by $\mathcal{O}$ the corresponding foliation of $M$ by $\phi$-orbits. Let $\tilde{M}$ be the universal cover of $M$ and let $\tilde{\mathcal{O}}$ be the lift of the foliation $\mathcal{O}$. The following is a fundamental result due to Barbot and Fenley: